# Hybrid Summary Statistics

### v1

@@ 1 Introduction
For many physics problems, such as large N-body and hydrodynamical solvers, forward simulations are expensive to generate, so networks tasked to compress these data for parameter inference must learn informative features using sparsely-sampled training sets, especially when large numbers of parameters are varied.
@@ 2 Formalism
[…]
@@
[…]
@@ 3 Experiments
In this initial experiment, we compare tree types of summaries, assuming a sufficiently large training set (10,000 samples): 1) Power spectrum only (11 $k$-bins). 2) Hybrid method: Power spectrum + two learned supplementary features (EPE Loss). 3) Hybrid method: Power spectrum + two learned supplementary features (CE Loss). One of the resulting inferences is shown in Figure 2. The results show that both approaches capture non-Gaussian information and improve inference performance, yielding similar outcomes. This suggests we can learn two additional parameters instead of new summaries, maintaining power spectrum interpretability while achieving near-optimal inference. The agreement also confirms both loss functions are effective, converging to the same results.
@@
Figure 3: Hybrid summaries are able to capture more non-Gaussian parameter information in settings with both high (left) and low (right) numbers of training simulations than a larger network trained with the same loss separately from and then concatenated to the $C_{\ell}$ statistic. When reducing the total available simulation volume all inferences suffer, but the hybrid statistics are most robust to this change, indicating that the MI objective improves capture of non-power spectrum features.
@@
We perform an ablation study to demonstrate the effectiveness of the hybrid statistics over neural-only methods and display results in Fig. 3. The embedding network in the hybrid setting is the lightweight CNN with symmetric kernels adapted from Makinen et al. (2024) to output 3 additional numbers. For comparison, we train a larger, more expressive CNN under the same EPE loss without access to the $C_{\ell}$ vector to output 3 numbers, and then for the density estimation step concatenate its outputs to the $C_{\ell}$s. We train the networks and density estimators first using all $5000$ simulations available (split into 70% train and 30% validation sets). We then reduce the total number of available simulations to $1000$ and further to $500$ and re-learn the embeddings and posteriors from scratch. When the simulation budget is reduced, all inferences suffer, but the hybrid statistic formalism encourages the smaller network to find non-Gaussian features in the dataset that are more robust to this change. We also note that the more expressive, separately-trained CNN inference degrades almost to the level of the $C_{\ell}$ contour in the lowest-data setting, even when concatenated to the $C_{\ell}$ vector in the density estimation step.
@@ Appendix A Mutual information maximization
This leads to a tractable loss function for maximizing mutual information, with $h(\theta)$ fixed, leading to the objective in Equation 2.
@@ B.2 Weak Lensing Simulations
Network details. For obtaining hybrid summaries the convolutional neural network with symmetric Multipole Kernels (MPK) was adapted from Makinen et al. (2024). For the large non-hybrid CNN we apply a $3\times 3$ kernel to embed the field into 16 filters, and then down-sample with stride-2 convolutions with output filters $[32,64,128]$. The network is then mean-pooled in the spatial axes and the flattened filters are passed to a dense network with a specified output size to be fed into the mutual information maximiser (here a mixture density network for the EPE loss). For both embedding networks we employ the smooth_leaky activation function from Makinen et al. (2024). For the EPE loss configuration, a mixture density network (MDN) with hidden layers of size $[70,70]$ and output layers for mixing coefficients, standard deviations, and means is employed to parameterise a four-component mixture.

### v2

@@ 1 Introduction
For many physics problems, such as large N-body and hydrodynamical solvers, forward simulations are expensive to generate, so networks tasked to compress these data for parameter inference must learn informative features using sparsely-sampled training sets, especially when large numbers of parameters are varied. Introducing informative priors in data feature (Battaglia et al., 2018; Ivanov et al., 2024, e.g.) or likelihood (Modi & Philcox, 2023) space can simplify the task of a neural network to learn an objective.
@@ 2 Formalism
In inference problems, we aim to infer quantities of interest $\theta$} from data d. In the physical sciences we often have domain knowledge which allows us to design a summary statistic of the data $\textbf{t}(\textbf{d})$ that capture some, but not all, of the information in a dataset. To enhance the information extraction, we aim to learn additional summaries $\textbf{s}(\textbf{d})$ that are complementary to t. We can formalise the extra information beyond t captured by s as the conditional mutual information $I(\textbf{s};{\theta}|\textbf{t})$, which measures how much the uncertainty about $\theta$} is reduced by knowing s given a static t. Evaluating the conditional mutual information can be difficult. However, using the chain rule of MI we can write
@@
where $\textbf{s}(\textbf{d})$ is a neural network and $q$ is a neural density estimator. Hoffmann & Onnela (2023) demonstrate that this loss unifies many information-theoretic losses into a stable objective. This mutual information objective can also be parameterised as a cross-entropy classification problem by employing the Jensen-Shannon divergence (Chen et al., 2021; Devon Hjelm et al., 2018; Nowozin et al., 2016):
@@ 3 Experiments
In this initial experiment, we compare tree types of summaries, assuming a sufficiently large training set (10,000 samples): 1) Power spectrum only (11 $k$-bins). 2) Hybrid method: Power spectrum + two learned supplementary features (EPE Loss). 3) Hybrid method: Power spectrum + two learned supplementary features (CE Loss). One of the resulting inferences is shown in Figure 2. The results show that both approaches capture non-Gaussian information and improve inference performance, yielding consistent posteriors. The slight difference in posterior recovery is likely due to differences between classifier and MDN network architectures for each loss, but each yielded similar convergence times in training. This suggests we can learn two additional parameters instead of new summaries, maintaining power spectrum interpretability while achieving near-optimal inference. The agreement also confirms both loss functions are effective, converging to the same results.
@@
Figure 3: Hybrid summaries are able to capture more non-Gaussian parameter information in settings with both high (left) and low (right) numbers of training simulations than summaries from a larger network trained with the same loss separately from and then concatenated to the $C_{\ell}$ statistic for the final inference in step 2. When reducing the total available simulation volume all inferences suffer, but the hybrid statistics are most robust to this change, indicating that the MI objective improves capture of non-power spectrum features.
@@
We perform an ablation study to demonstrate the effectiveness of the hybrid statistics over neural-only methods and display results in Fig. 3. The embedding network in the hybrid setting is the lightweight CNN with symmetric kernels adapted from Makinen et al. (2024) to output 3 additional numbers alongside the $C_{\ell}$s. For comparison, we train a larger, more expressive CNN embedding network under the same EPE loss without access to the $C_{\ell}$ vector to output 3 numbers, and then for the density estimation (step 2) concatenate its outputs to the $C_{\ell}$s (green contours; network (separate) + $C_{\ell}$). Step 2 probes the information content captured in the static network and $C_{\ell}$ summaries. We train the networks and density estimators from scratch first using all $5000$ simulations available (split into 70% train and 30% validation sets). We then reduce the total number of available simulations to $1000$ and further to $500$ and re-learn the embeddings and posteriors from scratch. When the simulation budget is reduced, all inferences suffer, but the hybrid statistic formalism encourages the smaller network to find non-Gaussian features in the dataset that are more robust to this change. We also note that the more expressive, separately-trained CNN inference degrades almost to the level of the $C_{\ell}$ contour in the lowest-data setting, even when concatenated to the $C_{\ell}$ vector in the density estimation step.
@@ 4 Conclusions & Outlook
5 Data Availability
All code, tutorials, and relevant simulations can be found at https://github.com/tlmakinen/hybridStats .
6 Acknowledgements
We thank Tao Jing, Justin Alsing, Tom Charnock, Niall Jeffrey, and David Spergel for conversations that inspired this work. This work was supported by the Simons Collaboration on “Learning the Universe”. TLM acknowledges the support of the Imperial College London President’s Scholarship and G-Research for conference travel costs. CS is supported by the National SKA Program of China (grant No. 2020SKA0110401) and NSFC (grant No. 11821303).
@@ Appendix A Mutual information maximization
This leads to a calculable loss function for maximizing mutual information, with $h(\theta)$ fixed, leading to the objective in Equation 2.
@@ B.2 Weak Lensing Simulations
Network details. For obtaining hybrid summaries the convolutional neural network with symmetric Multipole Kernels (MPK) was adapted from Makinen et al. (2024). The lightweight network is initialised without pretraining for this analysis and contains $1,615$ learnable parameters. For the large non-hybrid CNN we apply a $3\times 3$ kernel to embed the field into 16 filters, and then down-sample with stride-2 convolutions with output filters $[32,64,128]$. The network is then mean-pooled in the spatial axes and the flattened filters are passed to a dense network with a specified output size to be fed into the mutual information maximiser (here a mixture density network for the EPE loss). This results in $97,971$ learnable parameters. For both embedding networks we employ the smooth_leaky activation function from Makinen et al. (2024). For the EPE loss configuration, a mixture density network (MDN) with hidden layers of size $[70,70]$ and output layers for mixing coefficients, standard deviations, and means is employed to parameterise a four-component mixture.